# Twenty years of Nešetřil’s classification programme of Ramsey classes

### v3

@@ 6.2.1 Metric spaces
The construction of a completion of $G$ in Proposition 6.11 appears in various contexts. In Section 7.1 it will be useful that every automorphism of $G$ is also an automorphism of its completion. Another nice fact about it is that it is canonical; having canonical completions of free amalgams has turned out to be an important ingredient for understanding normal subgroups of automorphism groups of homogeneous structures, see e. g. [121, 289, 108, 202, 291, 292, 83, 181, 182, 183, 184, 68, 86, 173, 285, 185].
@@ 6.2.2 Equivalences
For $x\in C$, we will denote by $[x]$ its equivalence class in $\mathbf{C}$. Pick an arbitrary linear order $\prec$ of all equivalence classes in $\mathbf{C}$, and define a colouring $c\colon\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})\to 2$, such that, for an embedding $f\colon\mathbf{A}\to\mathbf{C}$, we put $c(f)=0$ if and only if $f(a_{0})]\prec[f(a_{1})]$, and $c(f)=1$ if and only if $f(a_{1})]\prec[f(a_{0})]$. Clearly, $c$ attains both colours on every embedding $\mathbf{B}\to\mathbf{C}$. ∎
When presented with a Ramsey problem, it is often a good strategy to look for bad colourings such as the one in Observation 6.14 as they often point in the direction of the desired expansion. In this concrete case, we know by Proposition 5.2 that we have to add an order, and consequently we will always have the structure $\mathbf{A}\in\mathcal{K}^{<}_{0}$ from the proof of Observation 6.14. Therefore, one needs to avoid the particular $\mathbf{B}$ from the proof. Note that the key property of $\mathbf{B}$ was that no matter how the equivalence classes are ordered, it contains both a copy of $\mathbf{A}$ where the order of vertices agrees with the order of their equivalence classes, as well as a copy of $\mathbf{A}$ where they disagree.
@@
Note that—as before—the existence of a homomorphism-embedding to $\mathbf{C}_{0}$ implies that the relation $_{\mathbf{C}}$ is acyclic. Consequently, we can fix a linear order $\prec_{I}$ on $I_{\mathbf{C}}$ extending $_{\mathbf{C}}$ as well as, for every $y\in I_{\mathbf{C}}$, a linear order $\prec_{y}$ on vertices $\{x\in O_{\mathbf{C}}:F_{\mathbf{C}}(x)=\{y\}\}$ extending $\leq_{\mathbf{C}}$. This allows us to define a binary relation $_{\mathbf{C}^{\prime}}$ on vertex set $C$ such that for every distinct $x,y\in C$ we have $x<_{\mathbf{C}^{\prime}}y$ if and only if one of the following holds:
@@
    $x,y\in O_{\mathbf{C}}$ and $F_{\mathbf{C}}(x)\prec_{I}F_{\mathbf{C}}(y)$, or
@@
    $x,y\in O_{\mathbf{C}}$, $F_{\mathbf{C}}(x)=F_{\mathbf{C}}(y)$, and $x\prec_{F_{\mathbf{C}}(x)}y$.
@@
This, together with $\prec_{I}$ extending $_{\mathbf{C}}$, implies that if there are vertices $x,y\in C$ with $x<_{\mathbf{C}}y$ and $y<_{\mathbf{C}^{\prime}}x$ then $x,y\in O_{\mathbf{C}}$. Since the orders $\prec_{z}$ extend $_{\mathbf{C}}$ inside the equivalence classes, it follows that $F_{\mathbf{C}}(x)\neq F_{\mathbf{C}}(y)$, and thus, by the construction of $_{\mathbf{C}^{\prime}}$, we know that $F_{\mathbf{C}}(y)\prec_{I}F_{\mathbf{C}}(x)$. Observe that the substructure of $\mathbf{C}$ with vertex set $\{x,y,F_{\mathbf{C}}(x),F_{\mathbf{C}}(y)\}$ is irreducible and hence embeds to $\mathbf{B}\in\mathcal{D}^{<}$. Consequently, since $x\leq_{\mathbf{C}}y$, we see that $F_{\mathbf{C}}(x)\leq_{\mathbf{C}}F_{\mathbf{C}}(y)$. This is in contradiction with $\prec_{I}$ extending $\leq_{\mathbf{C}}$.
@@ 6.2.3 The exceptions
Example 6 (Boolean algebras).
@@
Example 7 (Trees and $C$-relations).
@@ 6.3 A counterexample
What are the Ramsey expansions of $\mathcal{C}^{+}_{0}$ and $\mathcal{C}^{+}_{F}$? By the same argument as the one we discussed in Proposition 5.2 (showing that the automorphism group of every Ramsey structure fixes linear order of vertices), one can show that all extremely amenable subgroups of $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{0})$ and $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{F})$ must fix a 2-orientation. Since every vertex in $\mathbf{M}_{0}$ and $\mathbf{M}_{F}$ has infinite degree, there must be oriented paths of unbounded length [88, Theorem 3.7], which means that the expansions must be non-precompact.
This new example suggests that the traditional setup of KPT-correspondence can be extended. While neither $\mathcal{C}^{+}_{0}$ nor $\mathcal{C}^{+}_{F}$ have a precompact Ramsey expansion, it is still possible to obtain an optimal expansion in the sense that the corresponding extremely amenable subgroups $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{0})$ and $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{F})$ are maximal [88, Theorem 6.9]. See [282, 281] for further analysis of this example.
@@ 6.5 Some concepts from category theory
An example of a primal Ramsey theorem is Ramsey’s theorem itself which talks about finite linear orders with embeddings, and an example of a dual Ramsey theorem is the Graham–Rothschild theorem [111] which can (for the empty alphabet) be understood as a theorem about the category of finite linear orders with surjections that are rigid in the following sense: for every $x<y$ it holds that $\min(f^{-1}[x])<\min(f^{-1}[y])$. (The name rigid surjection was first used by Prömel and Voigt [245], though this view of the Graham–Rothschild theorem is due to Leeb [180].) Note that this is much stronger than the categorical dualisation of Ramsey’s theorem.
As far as we know, the term “dual Ramsey theorem” was first used in the 1980’s by Nešetřil and Rödl [220], and became established a few years later mainly thanks to the work of Carlson and Simpson [70]. Recently, Solecki [278] gave a general method for dualising Ramsey-type statements using the Galois connection: The objects which Ramsey-type theorems are concerned with typically contain natural partial orders (for the Ramsey theorem itself this is simply the linear order, but it can be a (model theoretic) tree in the case of the Milliken tree theorem [197], Deuber’s theorem for trees [72], or a related result of Leeb [112]). Given two partial orders $(S,\sqsubseteq_{S})$ and $(T,\sqsubseteq_{T})$ a pair $(f,e)$ is called a Galois connection if $f\colon T\to S$ and $e\colon S\to T$ are maps such that
@@
The partite construction is actually a method for producing structured Ramsey theorems over unstructured Ramsey categories with certain properties. This was already known to Nešetřil and Rödl, see also a recent paper of Junge [156] presenting the abstract partite construction in the categorical setting. In this survey we only focus on Prömel-indexed categories over finite linear orders, or in other words, classes of ordered structures, but an abstract framework, phrasing (as much as possible of) the constructions from Appendix A in an abstract categorical language, ideally implying analogues of Theorems 6.4 and 6.6 for Ramsey categories other than linear orders with embeddings, is currently being developed.
However, in some cases, the same Ramsey theorem can be viewed through multiple lenses. For example, the Graham–Rothschild theorem for rigid surjections is, as we have seen, an obviously unstructured Ramsey theorem. However, it turns out that it is equivalent to the Ramsey property of the class of all finite Boolean algebras with an antilexicographic order (see Example 6), which is a class of finite linearly ordered structures, and hence in the scope of Nešetřil’s classification programme. Despite this, it is visible from the proof of the Graham–Rothschild theorem that the dual unstructured setting is the natural setting for it (and it is thus, perhaps, not surprising that there is no known proof of the Ramsey property of the class of all finite Boolean algebras with an antilexicographic order using Theorem 6.7, see also Question 8.6).
Another example of this is the class of all finite (binary branching) $C$-relations with convex orders (see Example 7). Again, this is a class of finite relational structures with the Ramsey property. Nevertheless, these structures represent leaves of rooted binary trees with a lexicographic order. Moreover, the Ramsey property for this class is a consequence of the (finite variant of the) Milliken theorem [197] (see [39]), which is a generalization of the infinite Ramsey theorem to trees, if the Ramsey theorem is understood as a theorem about trees with no branchings. In particular, Milliken’s theorem is a primal unstructured Ramsey theorem, and thus, even though $C$-relations fall into the scope of Nešetřil’s classification programme, this explains why there is no known proof of the Ramsey property of the class of all finite (binary branching) $C$-relations with convex orders using Theorem 6.7.
@@
For more interactions of category theory and structural Ramsey theory see e. g. [107, 180, 215, 214, 218, 155, 224, 207, 175, 192, 75, 10, 12, 262, 50, 193, 195, 152, 156, 32, 279, 114].
@@ 7.1 Classification of EPPA classes
While EPPA and the Ramsey property may not seem related at first glance, there is an analogue of Observation 3.3 for EPPA:
@@ 7.1.1 Examples of classes with EPPA
Example 8 (Metric spaces).
@@
Since members of $\mathcal{D}$ are irreducible and automorphism-preserving completions are strong, we can use Lemma 6.13 to lift Hallbäck’s result to free orderings of $\mathcal{D}$ and $\mathcal{R}$, and apply Theorems 6.4 and 6.7. ∎
This answers a question of Hallbäck [117, Question A.6].
@@
Example 10 ($n$-partite tournaments).
@@
The other constructions in [87] or [124] also have a similar product structure and do not invoke Theorems 7.3 and 7.5. However, the structures handled in [165] needed a combination of the “ad hoc” product argument and Theorems 7.3 and 7.5 as they were metric-like. There, one needed the more general statements of Theorems 7.3 and 7.5 (where there is additionally a permutation group on the language and a morphism from $\mathbf{A}$ to $\mathbf{B}$ is a pair $(f,g)$ where $g$ is an element of the permutation group and $f$ is an embedding from $\mathbf{A}$ to $\mathbf{B}^{g}$ where the relations are permuted by $g$, see [128] for details). In general, it seems that a possible cookbook for proving EPPA might consists of combining such a product structure with Theorems 7.3 and 7.5.
@@ 7.2.1 Big Ramsey degrees
Theorem 7.10 (Laver, 1969, unpublished, see e.g. [288, Section 6.3]).
@@
Laver originally proved Theorem 7.10 by an argument described as “long and messy” by Devlin [74]. However, today it can be proved by a particularly elegant application of Milliken’s tree theorem [197], see [288, Section 6]. Milliken’s theorem itself builds on the well-known Halpern–Läuchli theorem [132], which was already known in 1969, but Laver was unaware of it.
@@
The odd tangent numbers are defined by the recurrence $t_{1}=1$ and $t_{k}=\sum_{\ell=1}^{k-1}\binom{2k-2}{2\ell-1}t_{l}\cdot t_{k-\ell}$, which happens to also count the number of certain special trees, called Devlin types [288, Section 6.3]. It is a well-known sequence (A000182 in the On-line Encyclopedia of Integer Sequences) starting by 1, 2, 16, 272, 7936, 353792, 22368256, 1903757312, 209865342976, …
@@ 7.2.3 Recent progress
For finite binary languages and finite sets $\mathcal{F}$, Zucker’s result directly implies an existential version of the Nešetřil–Rödl theorem (Theorem 2.6), that is, that there exists a Ramsey expansion, and with a careful analysis, one can even recover the exact expansion. Thus, Theorem 7.15 can be seen as an important step towards an infinite generalization of the Nešetřil–Rödl theorem. Again, both Dobrinen’s and Zucker’s proofs give upper bounds on the big Ramsey degrees only. A precise characterisation was subsequently obtained by Balko, Chodounský, Dobrinen, Hubička, Konečný, Vena, and Zucker [16] in a quite technically challenging analysis. See also [14, 131, 294] for a more compact presentation of the special cases of universal $\mathbf{K}_{3}$-free and $\mathbf{K}_{4}$-free graphs.
@@
This can be seen as a first step towards an finite generalization of Theorem 6.7. It is interesting that Theorem 7.17 does not demand that the number of forbidden cycles is finite. In particular, it also captures equivalences (interpreted as $\{N,E\}$-edge labelled graphs), ultrametric spaces with finitely many distances, $\Lambda$-ultrametric spaces [47], and more. This is in sharp contrast with Theorem 6.7 where one has to perform some gymnastics to pull these classes through. As a corollary, we get the following result for which we, at this point, do not have a general proof using Theorem 6.7 (although in all concrete cases it works, and we conjecture that there is also such a general proof):

### v5

@@ 6.2.1 Metric spaces
The construction of a completion of $G$ in Proposition 6.20 appears in various contexts. In Section 7.1 it will be useful that every automorphism of $G$ is also an automorphism of its completion. Another nice fact about it is that it is canonical; having canonical completions of free amalgams has turned out to be an important ingredient for understanding normal subgroups of automorphism groups of homogeneous structures, see e. g. [125, 293, 76, 112, 77, 78, 206, 295, 296, 87, 186, 187, 188, 189, 68, 90, 178, 289, 190].
@@ 6.2.2 Equivalences
For $x\in C$, we will denote by $[x]$ its equivalence class in $\mathbf{C}$. Pick an arbitrary linear order $\ll$ of all equivalence classes in $\mathbf{C}$, and define a colouring $c\colon\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})\to 2$, such that, for an embedding $f\colon\mathbf{A}\to\mathbf{C}$, we put $c(f)=0$ if and only if $f(a_{0})]\ll[f(a_{1})]$, and $c(f)=1$ if and only if $f(a_{1})]\ll[f(a_{0})]$. Clearly, $c$ attains both colours on every embedding $\mathbf{B}\to\mathbf{C}$. ∎
When presented with a Ramsey problem, it is often a good strategy to look for bad colourings such as the one in Observation 6.25 as they often point in the direction of the desired expansion. In this concrete case, we know by Proposition 5.9 that we have to add an order, and consequently we will always have the structure $\mathbf{A}\in\mathcal{K}^{<}_{0}$ from the proof of Observation 6.25. Therefore, one needs to avoid the particular $\mathbf{B}$ from the proof. Note that the key property of $\mathbf{B}$ was that no matter how the equivalence classes are ordered, it contains both a copy of $\mathbf{A}$ where the order of vertices agrees with the order of their equivalence classes, as well as a copy of $\mathbf{A}$ where they disagree.
@@
Note that—as before—the existence of a homomorphism-embedding to $\mathbf{C}_{0}$ implies that the relation $_{\mathbf{C}}$ is acyclic. Consequently, we can fix a linear order $\ll_{I}$ on $I_{\mathbf{C}}$ extending $_{\mathbf{C}}$ as well as, for every $y\in I_{\mathbf{C}}$, a linear order $\ll_{y}$ on vertices $\{x\in O_{\mathbf{C}}:F_{\mathbf{C}}(x)=\{y\}\}$ extending $\leq_{\mathbf{C}}$. This allows us to define a binary relation $_{\mathbf{C}^{\prime}}$ on vertex set $C$ such that for every distinct $x,y\in C$ we have $x<_{\mathbf{C}^{\prime}}y$ if and only if one of the following holds:
@@
    $x,y\in O_{\mathbf{C}}$ and $F_{\mathbf{C}}(x)\ll_{I}F_{\mathbf{C}}(y)$, or
@@
    $x,y\in O_{\mathbf{C}}$, $F_{\mathbf{C}}(x)=F_{\mathbf{C}}(y)$, and $x\ll_{F_{\mathbf{C}}(x)}y$.
@@
This, together with $\ll_{I}$ extending $_{\mathbf{C}}$, implies that if there are vertices $x,y\in C$ with $x<_{\mathbf{C}}y$ and $y<_{\mathbf{C}^{\prime}}x$ then $x,y\in O_{\mathbf{C}}$. Since the orders $\ll_{z}$ extend $_{\mathbf{C}}$ inside the equivalence classes, it follows that $F_{\mathbf{C}}(x)\neq F_{\mathbf{C}}(y)$, and thus, by the construction of $_{\mathbf{C}^{\prime}}$, we know that $F_{\mathbf{C}}(y)\ll_{I}F_{\mathbf{C}}(x)$. Observe that the substructure of $\mathbf{C}$ with vertex set $\{x,y,F_{\mathbf{C}}(x),F_{\mathbf{C}}(y)\}$ is irreducible and hence embeds to $\mathbf{B}\in\mathcal{D}^{<}$. Consequently, since $x\leq_{\mathbf{C}}y$, we see that $F_{\mathbf{C}}(x)\leq_{\mathbf{C}}F_{\mathbf{C}}(y)$. This is in contradiction with $\ll_{I}$ extending $\leq_{\mathbf{C}}$.
@@ 6.2.3 The exceptions
Example 6.29 (Boolean Algebras).
@@
Example 6.30 (Trees and $C$-relations).
@@ 6.3 A counterexample
What are the Ramsey expansions of $\mathcal{C}^{+}_{0}$ and $\mathcal{C}^{+}_{F}$? By the same argument as the one we discussed in Proposition 5.9 (showing that the automorphism group of every Ramsey structure fixes linear order of vertices), one can show that all extremely amenable subgroups of $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{0})$ and $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{F})$ must fix a 2-orientation. Since every vertex in $\mathbf{M}_{0}$ and $\mathbf{M}_{F}$ has infinite degree, there must be oriented paths of unbounded length [92, Theorem 3.7], which means that the expansions must be non-precompact.
This new example suggests that the traditional setup of KPT-correspondence can be extended. While neither $\mathcal{C}^{+}_{0}$ nor $\mathcal{C}^{+}_{F}$ have a precompact Ramsey expansion, it is still possible to obtain an optimal expansion in the sense that the corresponding extremely amenable subgroups $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{0})$ and $\mathop{\mathrm{Aut}}\nolimits(\mathbf{M}_{F})$ are maximal [92, Theorem 6.9]. See [285, 286] for further analysis of this example.
@@ 6.5 Some concepts from category theory
An example of a primal Ramsey theorem is Ramsey’s theorem itself which talks about finite linear orders with embeddings, and an example of a dual Ramsey theorem is the Graham–Rothschild theorem [115] which can (for the empty alphabet) be understood as a theorem about the category of finite linear orders with surjections that are rigid in the following sense: for every $x<y$ it holds that $\min(f^{-1}[x])<\min(f^{-1}[y])$. (The name rigid surjection was first used by Prömel and Voigt [249], though this view of the Graham–Rothschild theorem is due to Leeb [185].) Note that this is much stronger than the categorical dualization of Ramsey’s theorem.
As far as we know, the term “dual Ramsey theorem” was first used in the 1980’s by Nešetřil and Rödl [224], and became established a few years later mainly thanks to the work of Carlson and Simpson [70]. Recently, Solecki [282] gave a general method for dualizing Ramsey-type statements using the Galois connection: The objects which Ramsey-type theorems are concerned with typically contain natural partial orders (for the Ramsey theorem itself this is simply the linear order, but it can be a (model theoretic) tree in the case of the Milliken tree theorem [201], Deuber’s theorem for trees [72], or a related result of Leeb [116]). Given two partial orders $(S,\sqsubseteq_{S})$ and $(T,\sqsubseteq_{T})$ a pair $(f,e)$ is called a Galois connection if $f\colon T\to S$ and $e\colon S\to T$ are maps such that
@@
The partite construction is actually a method for producing structured Ramsey theorems over unstructured Ramsey categories with certain properties. This was already known to Nešetřil and Rödl, see also a recent paper of Junge [161] presenting the abstract partite construction in the categorical setting. In this survey we only focus on Prömel-indexed categories over finite linear orders, or in other words, classes of ordered structures, but an abstract framework, phrasing (as much as possible of) the constructions from Appendix A in an abstract categorical language, ideally implying analogues of Theorems 6.5 and 6.11 for Ramsey categories other than linear orders with embeddings, is currently being developed.
However, in some cases, the same Ramsey theorem can be viewed through multiple lenses. For example, the Graham–Rothschild theorem for rigid surjections is, as we have seen, an obviously unstructured Ramsey theorem. However, it turns out that it is equivalent to the Ramsey property of the class of all finite Boolean algebras with an antilexicographic order (see Example 6.29), which is a class of finite linearly ordered structures, and hence in the scope of Nešetřil’s classification programme. Despite this, it is visible from the proof of the Graham–Rothschild theorem that the dual unstructured setting is the natural setting for it (and it is thus, perhaps, not surprising that there is no known proof of the Ramsey property of the class of all finite Boolean algebras with an antilexicographic order using Theorem 6.15, see also Question 8.6).
Another example of this is the class of all finite (binary branching) $C$-relations with convex orders (see Example 6.30). Again, this is a class of finite relational structures with the Ramsey property. Nevertheless, these structures represent leaves of rooted binary trees with a lexicographic order. Moreover, the Ramsey property for this class is a consequence of the (finite variant of the) Milliken theorem [201] (see [39]), which is a generalization of the infinite Ramsey theorem to trees, if the Ramsey theorem is understood as a theorem about trees with no branchings. In particular, Milliken’s theorem is a primal unstructured Ramsey theorem, and thus, even though $C$-relations fall into the scope of Nešetřil’s classification programme, this explains why there is no known proof of the Ramsey property of the class of all finite (binary branching) $C$-relations with convex orders using Theorem 6.15.
@@
For more interactions of category theory and structural Ramsey theory see e. g. [111, 185, 219, 218, 222, 160, 228, 211, 180, 197, 79, 10, 12, 266, 50, 199, 157, 161, 32, 283, 118].
@@ 7.1 Classification of EPPA classes
While EPPA and the Ramsey property may not seem related at first glance, there is an analogue of Observation 3.6 for EPPA:
@@ 7.1.1 Examples of classes with EPPA
Example 7.8 (Metric spaces).
@@
Since members of $\mathcal{D}$ are irreducible and automorphism-preserving completions are strong, we can use Lemma 6.23 to lift Hallbäck’s result to free orderings of $\mathcal{D}$ and $\mathcal{R}$, and apply Theorems 6.5 and 6.15. ∎
This answers a question of Hallbäck [121, Question A.6].
@@
Example 7.12 ($n$-partite tournaments).
@@
The other constructions in [91] or [128] also have a similar product structure and do not invoke Theorems 7.5 and 7.7. However, the structures handled in [170] needed a combination of the “ad hoc” product argument and Theorems 7.5 and 7.7 as they were metric-like. There, one needed the more general statements of Theorems 7.5 and 7.7 (where there is additionally a permutation group on the language and a morphism from $\mathbf{A}$ to $\mathbf{B}$ is a pair $(f,g)$ where $g$ is an element of the permutation group and $f$ is an embedding from $\mathbf{A}$ to $\mathbf{B}^{g}$ where the relations are permuted by $g$, see [133] for details). In general, it seems that a possible cookbook for proving EPPA might consists of combining such a product structure with Theorems 7.5 and 7.7.
@@ 7.2.1 Big Ramsey degrees
Theorem 7.16 (Laver, 1969, Unpublished, See e.g. [292, Section 6.3]).
@@
Laver originally proved Theorem 7.16 by an argument described as “long and messy” by Devlin [74]. However, today it can be proved by a particularly elegant application of Milliken’s tree theorem [201], see [292, Section 6]. Milliken’s theorem itself builds on the well-known Halpern–Läuchli theorem [137], which was already known in 1969, but Laver was unaware of it.
@@
The odd tangent numbers are defined by the recurrence $t_{1}=1$ and $t_{k}=\sum_{\ell=1}^{k-1}\binom{2k-2}{2\ell-1}t_{l}\cdot t_{k-\ell}$, which happens to also count the number of certain special trees, called Devlin types [292, Section 6.3]. It is a well-known sequence (A000182 in the On-line Encyclopedia of Integer Sequences) starting by 1, 2, 16, 272, 7936, 353792, 22368256, 1903757312, 209865342976, …
@@ 7.2.3 Recent progress
For finite binary languages and finite sets $\mathcal{F}$, Zucker’s result directly implies an existential version of the Nešetřil–Rödl theorem (Theorem 2.9), that is, that there exists a Ramsey expansion, and with a careful analysis, one can even recover the exact expansion. Thus, Theorem 7.22 can be seen as an important step towards an infinite generalization of the Nešetřil–Rödl theorem. Again, both Dobrinen’s and Zucker’s proofs give upper bounds on the big Ramsey degrees only. A precise characterization was subsequently obtained by Balko, Chodounský, Dobrinen, Hubička, Konečný, Vena, and Zucker [15] in a quite technically challenging analysis. See also [16, 136, 298] for a more compact presentation of the special cases of universal $\mathbf{K}_{3}$-free and $\mathbf{K}_{4}$-free graphs.
@@
This can be seen as a first step towards an infinite generalization of Theorem 6.15. It is interesting that Theorem 7.24 does not demand that the number of forbidden cycles is finite. In particular, it also captures equivalences (interpreted as $\{N,E\}$-edge labelled graphs), ultrametric spaces with finitely many distances, $\Lambda$-ultrametric spaces [46], and more. This is in sharp contrast with Theorem 6.15 where one has to perform some gymnastics to pull these classes through. As a corollary, we get the following result for which we, at this point, do not have a general proof using Theorem 6.15 (although in all concrete cases it works, and we conjecture that there is also such a general proof):